# AQM: A Refresh of the Abstract Qubit Model for Quantum Co-design

### v1

@@ Abstract
Qubits are the fundamental building blocks of quantum information science and applications, whose concept is widely utilized in both quantum physics and quantum computation. While the significance of qubits and their implementation in physical devices have been extensively examined, we posit that now is the right time to revisit this understanding. In this paper, we introduce an abstract qubit model (AQM), offering a mathematical framework for higher-level algorithms and applications, and setting forth criteria for lower-level physical devices to enable quantum computation. We first provide a comprehensive definition of “qubits”, regarded as the foundational premise for quantum algorithms (bottom-up support), and examine their requisites for devices (top-down demand). We then investigate the feasibility of moderating specific requirements, thereby broadening device support while considering techniques that tradeoff extra costs to counterbalance this moderation. Lastly, we delve into the quantum applications that only require incomplete qubits, and discuss the physical systems having restricted AQM support but are still useful in quantum applications. AQM may serve as an intermediate interface between quantum algorithms and devices, facilitating quantum algorithm-device co-design.
I Introduction
@@
Quantum computing (QC) holds the potential for significant acceleration of classically challenging problems Grover (1996); Shor (1994, 1997); Simon (1997), while quantum information processing (QIP) facilitates secure communication underpinned by quantum mechanics Bennett and Brassard (1984); Xu et al. (2020); Portmann and Renner (2022). Along with the rapid advancement of quantum technology is the growing selection of physical systems promising for QC and QIP. Despite these devices being rooted in diverse materials and fabrication processes, certain essential functionalities exist, such as initialization, control, and readout of quantum information, that every device is deemed to offer. These functionalities represent the commonality, or the criteria, necessary for a device to be capable of engaging in QC and QIP.
@@
In order to examine whether a physical system can be a promising candidate for QC and QIP, DiVincenzo in his seminal work DiVincenzo (2000) proposed the well-known criteria. One of the key concepts, which will also be the main focus of our paper, is the concept of qubit.
@@
As the fundamental building blocks of gate-based quantum computers, the abstraction of qubits is necessary to build a model for quantum computers and support the upper-stack design. As described by Nielsen and Chuang, a qubit is defined as a mathematical object, each of which has a state that is a unit vector in a two-dimensional complex vector space Nielsen and Chuang (2010). Looking from the bottom of the QC design stack, this criteria implies also the requirements placed on the physical platforms. The physical platform should support such a system, abstracted as a qubit. To better understand such a qubit in realistic physical systems, especially when the physical degrees of freedom are more than a qubit, in Ref. Viola et al. (2001), Viola et al. proposed that: in addition to the operator algebra, the physical system should also support (1) universal control, (2) initialization, and (3) read-out capabilities to build an operational qubit.
@@
Despite classical, it has been more than 20 years since DiVincenzo’s qubit criteria. During the recent development of QC and QIP, new opportunities and challenges have arisen. Firstly, although a large integration of physical qubits has been demonstrated Gambetta (2022); Bluvstein et al. (2024); Gambetta (2023), a perfect logical qubit has not been built yet. We still need to take advantage of the hardware features to maximize the benefits in QC. Second, some recent quantum algorithms do not require a complete AQM to support. The rigorous AQM has the potential to be released. Beyond the commonly used quantum circuit model, various universal computation models exist, such as measurement-based quantum computation (MBQC) Raussendorf and Briegel (2001); Raussendorf et al. (2003); Browne and Rudolph (2005); Briegel et al. (2009) and quantum random walk Kadian et al. (2021); Venegas-Andraca (2012); Childs (2009); Singh et al. (2021). These models have unique requirements for physical systems. Lastly, with the recent discussion of quantum co-design Shi et al. (2020); Tomesh and Martonosi (2021), partially supporting AQM can benefit both quantum algorithm designs and hardware development. How to leverage the AQM for algorithm-device co-design is still unclear. These emerging topics call for a reconsideration of the AQM.
@@
In this paper, we address these challenges by revisiting the AQM. We begin by presenting a comprehensive definition of qubits, essential for quantum algorithms (bottom-up support), and analyze their device requirements (top-down demand). We explore the opportunities of relaxing specific requirements to expand device support, considering techniques that trade off extra costs to compensate for the relaxation. Finally, we examine quantum applications that only need incomplete qubits and discuss physical systems with limited AQM support yet remain valuable in quantum applications.
@@
This paper is organized as follows. In Sec. II, we define the abstracted qubit model. In Sec. III, we discuss how the physical requirements can be released and compensated. In Sec. IV, we focus on releasing the complete AQM from both the upper and lower stacks opportunities. Finally, we conclude in Sec. V.
@@
Figure 1: The abstracted qubit model in the design stacks of quantum information processing. The abstracted qubit model contains the mathematical description of qubits, which can be used in the upper stacks, while it places physical requirements that need to be fulfilled by the lower stacks.
II Abstract model of qubits
@@
We present the definition of qubits and construct the AQM. The AQM design stack is shown in Fig. 1. Specifically, we consider how qubits can support the upper stacks of QC and QIP systems by defining the mathematical description of the states and the operations on qubits in Sec. II.1. Then, we discuss the requirements of AQM placed on the physical systems in Sec. II.2.
II.1 Mathematical description of qubits
@@
  • State of qubits: The state of a qubit can be represented as a complex two-dimensional vector with the unit norm, same as Ref. Nielsen and Chuang (2010). The state of $n$ qubits is a complex vector with dimension $2^{n}$. Assuming the state of qubit A is |$\ket{A}$, while the state of qubit B is |$\ket{B}$, the state of both qubits is |$\ket{A}\otimes\ket{B}$, where $\otimes$ represent the Kronecker product of two vectors.
@@
Qubit vs Qudit: With the possibility of data compression and the advantages of simulating the dynamics with other symmetry groups, qudits become an attractive generalization and alternative to qubits Wang et al. (2020); Chi et al. (2022); Liu et al. (2023a); González-Cuadra et al. (2022). A qudit is a system with $d$ states, where $d$ can be more than $2$. The mathematical definition discussed above can be easily generalized to qudit systems by redefining the dimensionality of a single qudit to $d$, while the operation matrix dimensions are $d\times d$. In the following of our discussion, we mainly focus on qubits, while the generalization to qudit systems is straightforward as a simple linear mapping.
II.2 Physical requirements for constructing qubits
Recall DiVincenzo’s five criteria DiVincenzo (2000) for examining physical systems for QC:
@@ II.2 Physical requirements for constructing qubits
Subsequent works refine qubit requirements and explore building qubits for QC and QIP Viola et al. (2001); Duckering (2022). Here we focus on DiVincenzo’s criteria because of the wide adoption in QC. The ideal qubit model contains:
@@
  • Rule 2: Operations. The capability of performing operations on qubits requires the physical system that makes these $n$ qubits to be completely controllable Dong and Petersen (2010); D’Alessandro (2021). Obviously, with the complete controllability of the physical systems, universal quantum computing can be supported.
@@
  • Rule 5: Readout. In addition, as projective measurements and post-selection results can be well described by the measurement formalism Nielsen and Chuang (2010), with the above measurement operation support, the qubit states can be readout into classical information. We also implicitly require that the measurement is error-free, meaning that the extracted classical information is accurate. The initialization of the qubits to a known state (noted as |$\ket{\psi_{0}}$) can also be realized through this measurement operation by defining the measurement operator set $\mathbf{M}=\{\outerproduct{\psi_{0}}{j}\}$ | } for all basis states of the qubit systems |$\ket{j}$.
III Incomplete support of the AQM from physical systems
Table 1: Comparison of the ideal AQM and the incomplete AQM. We also consider the techniques to compensate for the imperfection and their corresponding costs.
Ideal model
Restricted model
Compensation techniques
Cost
Two quantum states
Two quantum states
Infinitely long coherence time
Finite but long coherence time
Quantum Error Detection and Correction
More physical qubits and more gate and measurement operations
Quantum error mitigation
Extra quantum algorithm implementations, and post-processing loads.
Dynamical decoupling
Extra control complexities
No quantum operation errors
gate imperfection are allowed
Quantum Error Detection and Correction
More physical qubits and more gate and measurement operations
Directly support any unitary operations.
A discrete universal gate set is available.
Gate decomposition
More gate operations, which require more operation time. Needs more qubits if the gate set is computationally universal,
All-to-all connectivity.
All qubits can be connected.
Gate routing
More physical operations and extra times.
General measurements are accessible.
Projective measurements can be performed, even indirectly.
Measurements with auxiliary qubits
Extra qubits and quantum gates to construct a general measurement.
Routing for measurements
Swapping the quantum states to the qubits requires extra gate operations and time.
Figure 2: The incomplete support of the AQM from the physical systems. With compensation techniques, the physical systems can support the complete AQM with extra costs on time and resources.
@@ III Incomplete support of the AQM from physical systems
Relaxing qubit states: The qubit should have two quantum states.
@@
The existence of two quantum states for a qubit is necessary. One commonly seen example is in trapped ion or neutral atom systems, where the quantum levels of the ions and atoms are used to encode quantum information. Another example is to use the presence or absence of a particle (photons or electrons, etc.) in a single mode as the two quantum states of the qubits. We will discuss it in more detail in Sec. IV.2.3. There may be more than two quantum states in a certain physical system Viola et al. (2001). Other than simply picking up two distinct states from all available quantum states, the quantum states need to be controllable (see the following discussion about Operation).
@@
On the other side, some physical platforms may not have enough quantum states that fulfill even the minimum requirements of a single qubit, e.g., the space formed by the two states can be addressed, and the operations on the qubits should map the state space back to itself. Similar constraints apply to qudit constructions. When a single physical system is not sufficient to provide such a state space, however, the composite system of multiple physical systems can be utilized to form a qubit, e.g., the presence or absence of a single electron in a quantum dot cannot be encoded as qubit state due to the lack of coherent manipulation techniques in the Hilbert space spanned by these two states DiVincenzo (2000). In contrast, the presence or absence of a microwave photon can be encoded as a qubit with the help of a superconducting qubit Mariantoni et al. (2011a, b); Besse et al. (2020); Reuer et al. (2022); Gao et al. (2019).
@@
We note that this requirement is weaker than the requirements of single-qubit completely controllable (Rule 2) Dong and Petersen (2010); D’Alessandro (2021). Despite incomplete control over the physical system, if a set of computational universal gates can be executed, and the qubit state can be manipulated to a specific state (e.g., |$\ket{0}$), the construction of a qubit is possible. The arbitrary state of the qubit Hilbert space can be approximately approached with the extra cost of quantum gate operations (see the discussion of Operations).
@@
The quantum information encoded in qubit Hilbert space can be transmitted via various physical systems. In measurement-based quantum computing (MBQC) Raussendorf and Briegel (2001); Raussendorf et al. (2003); Raussendorf and Harrington (2007); Briegel et al. (2009), for instance, information is ”teleported” to connected physical qubits in resource states, with the physical qubits being destroyed during computation, especially in photonic systems.
@@
Relaxing operations: The qubit should support a set of computational universal gates.
@@
It is quite costly to require any unitary matrices on an arbitrary number of qubits to be directly implementable as different operations. What is the minimum requirement on the set of supported operations to allow us, at least approximate a given unitary matrix on a number of qubits? This question motivates the research of the universal gate set of QC. It turns out that a discrete set of gates (unitary matrices) on a small number of qubits is enough ($n\leq 3$), e.g., the set of Hadamard gate, CNOT gate, phase gate, and $\pi/8$ gate is universal Nielsen and Chuang (2010).
@@
However, releasing this requirement also brings extra cost, which has two aspects. One is from the gate decomposition, which decomposes the required unitary matrix into the directly supported gate operations in the gate set. The gate decomposition requires an increasing number of quantum gates applied to the qubits. For example, the Solovay-Kitaev theorem showed that approximating an arbitrary unitary matrix in $SU(d)$ using a discrete set of gates that generate a dense subgroup of $SU(d)$ requires $O(log^{c}(1/\epsilon))$ gates, where $c$ is a constant and $\epsilon$ is the accuracy Nielsen and Chuang (2010); Dawson and Nielsen (2005). The second aspect is that the different universal gate sets supported by the physical system can cost extra quantum gates and qubits for the same computation, as highlighted by the difference between ‘computational universal’ and ‘strictly universal’ Aharonov (2003). One example is the gate set that consists of the Toffoli gate and the Hadamard gate. This gate set is universal Aharonov (2003); Shi (2002), but a general unitary on a $n$-qubit system cannot be decomposed with arbitrary accuracy using just Toffoli and Hadamard gates. Instead, the computation process of that unitary operation can be simulated within an arbitrarily small error Aharonov (2003). However, it costs polynomially many more qubits and gates. Similar discussions about universality can be found in qudit systems Brylinski and Brylinski (2001); Wang et al. (2020). We stress that the cost of decomposing target unitaries on the qubit system should be manageable, e.g., not using exponentially more time and physical qubits.
@@
Nevertheless, there are other QC models other than the quantum circuit models. For example, in the MBQC scheme Raussendorf and Briegel (2001); Raussendorf et al. (2003); Raussendorf and Harrington (2007); Briegel et al. (2009), universal computing is driven by the measurements on a prepared entangled state. In this model, the computation can still be mapped back to the circuit model, where the flow of quantum information can be treated as qubits that carry the quantum information. The physical systems that support MBQC can still be treated as supporting full universal control, though indirectly with the help of measurements and pre-existed quantum entanglements.
Relaxing connectivity: The qubits needs to be connected.
@@
For a QC system supporting a universal gate set applicable up to $n$ qubits, a prerequisite for using these qubits in general QC algorithms is that any subset of $n$ qubits must be capable of performing gates from the universal set. Note that all-to-all connectivity is not necessary. For example, if the universal gate set contains CNOT, Hadamard, phase and $\pi/8$ gate Nielsen and Chuang (2010), this is equivalent to requiring that all qubits can perform the Hadamard, phase and $\pi/8$ gates, and any two qubits can find a path between in their connection graph. Even if two qubits are not directly connected, a CNOT gate between them can be implemented using SWAP gates along the path, decomposable into three CNOTs. This less stringent connectivity requirement incurs additional costs compared to all-to-all connectivity.
@@
In the new paradigm of distributed QC models, the “connection” between the two qubits we discussed above is not necessarily made of physical contacts, or direct coupling between two physical qubits. Instead, quantum communication between two physically remote qubits enables remote gate operations Cuomo et al. (2020); Cacciapuoti et al. (2020), which relies on a reliable channel to transmit flying qubits from one end to the other, or a reliable source of entangled resource states (like Bell states) to be used for gate teleportation Gottesman and Chuang (1999); Huang et al. (2004); Wan et al. (2019); Chou et al. (2018); Nielsen and Chuang (2010). This should also comply with DiVincezo’s criteria for quantum communication (criteria 6 and 7, see Sec. II.2).
@@
Relaxing coherence: The qubit should have finite but long enough coherence time.
1. relaxing to the coherence time enough for gate operation, 2. using QEC to extend logical qubit coherence.
@@
Quantum coherence is critical for quantum computing. However, physical systems can hardly be completely isolated. The interaction with the external environment causes decoherence of the quantum systems. The coherence time gives a time scale for preserving the quantum coherence. These unwanted interactions also introduce decoherence errors to the quantum gates applied to the systems. Therefore, the coherence time should be long enough for the gate operations time to perform a meaningful QC on a physical system. This motivates the development of noisy intermediate-scale quantum (NISQ) algorithms, e.g., variational quantum algorithms Cerezo et al. (2021)
@@
Another strategy is to suppress the noise and extend the coherence time of the logical qubits by quantum error correction (QEC) codes Nielsen and Chuang (2010); Calderbank and Shor (1996). By combining multiple physical qubits into a logical qubit, QEC codes use entangled states as qubit states to detect and correct errors. Increasing the code distance, achieved by using more physical qubits per logical qubit, allows for suppressing logical error rates as long as noise-induced error rates remain below the code threshold. What’s more, QEC not only extends the coherence time of logical qubits but also suppresses noises from other sources, e.g., imperfect control during gate implementation and cross-talk between adjacent qubits, etc.
@@
Relaxing readout: The qubit should have a way of performing measurements on qubit states.
@@
Furthermore, even the direct projective measurements on the physical systems can be released if we consider heterogeneous QC architectures, where the qubits used in different functional units inside the QC units Stein et al. (2023a, b); Liu et al. (2023b) can be different. The role of the qubit used in the quantum processing units and quantum memory units is to process and store quantum information, which means direct measurements on them are not necessary. We can think of using another physical system, that supports qubit measurement functionalities and can interact with the qubits in the quantum memory modules to extract the stored quantum states. Therefore, when the measurement of the quantum information is needed, it is performed on the second type of qubits by swapping the information out from the computing qubits. A more detailed discussion is in Sec. IV.1.5.
@@
Figure 3: The algorithm-device co-design. Quantum algorithms and applications may only have partial demand on the AQM, while the physical systems can only have partial support of the AQM. The co-design of quantum algorithms and physical systems can be beneficial.
IV AQM for Algorithm-Device codesign
In this section, we discuss the algorithm-device co-design opportunities demonstrated by Fig. 3. Specifically, a quantum application may not require the complete support of the AQM, and hence a physical system that fulfills its requirements can still be useful, even without building a qubit, and vice versa. Therefore, the algorithm-device codesign opportunities by partially breaking the abstraction layer of qubits can be beneficial for the current stage of QC system design. Specifically, in Sec. IV.1, we discuss the algorithm and applications that can be supported without the complete AQM, while in Sec. IV.2, we discuss different physical systems supporting the AQM or quantum applications. In Sec. IV.3, we discuss co-design opportunities of the physical systems.

### v2

# AQM: A Refresh of the Abstract Qubit Model for Quantum Computing Co-design

@@ Abstract
Qubits are the fundamental building blocks of quantum information science and applications, whose concept is widely utilized in both quantum physics and quantum computation. While the significance of qubits and their implementation in physical devices have been extensively examined, now is the right time to revisit this understanding. In this paper, we introduce an abstract qubit model (AQM), offering a mathematical framework for higher-level algorithms and applications, and setting forth criteria for lower-level physical devices to enable quantum computation. We first provide a comprehensive definition of “qubits”, regarded as the foundational principle for quantum computing algorithms (bottom-up support), and examine their requisites for devices (top-down demand). We then investigate the feasibility of relaxing specific requirements, thereby broadening device support while considering techniques that tradeoff extra costs to counterbalance this relaxation. Lastly, we delve into the quantum applications that only require partial support of “qubits”, and discuss the physical systems with limited support of the AQM but remain valuable in quantum applications. AQM may serve as an intermediate interface between quantum algorithms and devices, facilitating quantum algorithm-device co-design.
1 Introduction
@@
Quantum computing (QC) holds the potential for significant acceleration of classically challenging problems [1, 2, 3, 4], while quantum information processing (QIP) facilitates secure communication underpinned by quantum mechanics [5, 6, 7]. Along with the rapid advancement of quantum technology is the growing selection of physical systems promising for QC and QIP. Despite these devices being rooted in diverse materials and fabrication processes, certain essential functionalities exist, such as initialization, control, and readout of quantum information, that every device is deemed to offer. These functionalities represent the commonality, or the criteria, necessary for a device to be capable of engaging in QC and QIP.
@@
In order to examine whether a physical system can be a promising candidate for QC and QIP, DiVincenzo in his seminal work [8] proposed the well-known criteria. One of the key concepts, which will also be the main focus of our paper, is the concept of qubits.
@@
As the fundamental building blocks of gate-based quantum computers, the abstraction of qubits is necessary to build a model for quantum computers and support the upper-stack design. As described by Nielsen and Chuang, a qubit is defined as a mathematical object, each of which has a state that is a unit vector in a two-dimensional complex vector space [9]. Looking from the bottom of the QC design stack, this criteria implies also the requirements placed on the physical platforms. The physical platform should support such a system, abstracted as a qubit. To better understand such a qubit in realistic physical systems, especially when the physical degrees of freedom are more than a qubit required, in Ref. [10], Viola et al. proposed that: in addition to the operator algebra, the physical system should also support (1) universal control, (2) initialization, and (3) read-out capabilities to build an operational qubit.
@@
It has been more than 20 years since DiVincenzo’s qubit criteria. During the recent development of QC and QIP, new opportunities and challenges have arisen. Firstly, although a large integration of physical qubits has been demonstrated [11, 12, 13], a perfect logical qubit has not been built yet. We still need to take advantage of hardware features to maximally improve the QC development. Second, some recent quantum algorithms can still function without the full support of the AQM, and hence the rigorous AQM has the potential to be further relaxed. Beyond the commonly used quantum circuit model, various universal computation models exist, such as measurement-based quantum computation (MBQC) [14, 15, 16, 17] and quantum random walk [18, 19, 20, 21]. These models have unique requirements for physical systems. Lastly, with the recent progress of quantum co-design [22, 23], partially supporting AQM can provide benefits to both quantum algorithm design and hardware development. How to leverage the AQM for algorithm-device co-design is still unclear. These emerging topics call for a reconsideration of the AQM.
@@
In this paper, we address these challenges by revisiting the AQM. We begin by presenting a comprehensive definition of qubits, essential for quantum algorithms (bottom-up support), and analyze their device requirements (top-down demand). We explore the opportunities of relaxing specific requirements to expand device support, considering techniques that trade off extra costs to compensate for the relaxation. Finally, we examine quantum applications that only need partial AQM functionalities and discuss physical systems with limited AQM support yet remain valuable in quantum applications.
@@
This paper is organized as follows. In Sec. 2, we define the abstracted qubit model. In Sec. 3, we discuss how the physical requirements can be relaxed and compensated. In Sec. 4, we focus on relaxing the complete AQM from both the upper and lower stacks opportunities. Finally, we conclude in Sec. 5.
@@
Figure 1: The abstracted qubit model in the design stacks of quantum information processing. The abstracted qubit model contains the mathematical description of qubits, which can be used in the upper stacks, while it places physical requirements that need to be fulfilled by the lower stacks. The complete AQM can assure the function execution in quantum algorithms (the red arrow), while it also places demands on the physical systems that construct qubits (the blue arrow).
2 Abstract model of qubits
@@
We present the definition of qubits and construct the AQM. The AQM design stack is shown in Fig. 1. Specifically, we consider how qubits can support the upper stacks of QC and QIP systems by defining the mathematical description of the states and the operations on qubits in Sec. 2.1. Then, we discuss the requirements of AQM placed on the physical systems in Sec. 2.2.
2.1 Mathematical description of qubits
@@
  • State of qubits: The state of a qubit can be represented as a complex two-dimensional vector with the unit norm, same as Ref. [9]. The state of $n$ qubits is a complex vector with dimension $2^{n}$. Assuming the state of qubit A is |$\ket{A}$, while the state of qubit B is |$\ket{B}$, the state of both qubits is |$\ket{A}\otimes\ket{B}$, where $\otimes$ represent the Kronecker product of two vectors.
@@
Qubit vs Qudit: With the possibility of data compression and the advantages of simulating the dynamics with other symmetry groups, qudits become an attractive generalization and alternative to qubits [24, 25, 26, 27]. A qudit is a system with $d$ states, where $d$ can be more than $2$. The mathematical definition discussed above can be easily generalized to qudit systems by redefining the dimensionality of a single qudit to $d$, while the operation matrix dimensions are $d\times d$. In the following of our discussion, we mainly focus on qubits, while the generalization to qudit systems is a straightforward simple linear mapping.
2.2 Physical requirements for constructing qubits
Recall DiVincenzo’s five criteria [8] for examining physical systems for QC:
@@ 2.2 Physical requirements for constructing qubits
Subsequent works refine qubit requirements and explore building qubits for QC and QIP [10, 28]. Here we focus on DiVincenzo’s criteria because of the wide adoption in QC. The ideal qubit model contains:
@@
  • Rule 2: Operations. The capability of performing operations on qubits requires the physical system that makes these $n$ qubits to be completely controllable [29, 30]. Obviously, with the complete controllability of the physical systems, universal quantum computing can be supported.
@@
  • Rule 5: Readout. In addition, as projective measurements and post-selection results can be well described by the measurement formalism [9], with the above measurement operation support, the qubit states can be readout into classical information. We also implicitly require that the measurement is error-free, meaning that the extracted classical information is accurate. The initialization of the qubits to a known state (noted as |$\ket{\psi_{0}}$) can also be realized through this measurement operation by defining the measurement operator set $\mathbf{M}=\{\outerproduct{\psi_{0}}{j}\}$ | } for all basis states of the qubit systems |$\ket{j}$.
3 Partial support of the AQM from physical systems
Table 1: Comparison of the ideal AQM and the partial AQM. We also consider the techniques to compensate for the imperfection and their corresponding costs.
Ideal model Partial model Compensation techniques Cost
Two quantum states Two quantum states
Infinitely long coherence time Finite but relatively long coherence time Quantum Error Detection and Correction More physical qubits and more gate and measurement operations
Quantum error mitigation Extra quantum algorithm implementations, and post-processing loads.
Dynamical decoupling Extra control complexities
No quantum operation errors Gate imperfection are allowed Quantum Error Detection and Correction More physical qubits, gates, and measurement operations
Directly support any unitary operations. A discrete universal gate set is available. Gate decomposition Increased gate operations, which require more operation time. Needs more qubits if the gate set is computationally universal.
All-to-all connectivity. All qubits can be connected. Gate routing More physical operations and extra times.
General measurements are accessible. Projective measurements can be performed, even indirectly. Measurements with auxiliary qubits Extra qubits and quantum gates to construct a general measurement.
Routing for measurements Swapping the quantum states to the qubits requires extra gate operations and time.
Figure 2: The partial support of the AQM from the physical systems. With compensation techniques, the physical systems can support the complete AQM with extra costs on time and resources. The physical systems can partially support the AQM (the blue arrow), while the AQM can assure the function implementation in quantum algorithms (the red arrow). Compensation techniques can be applied to support the complete AQM (the gray arrow).
@@ 3 Partial support of the AQM from physical systems
a. Relaxing the qubit state requirement: The qubit should have two quantum states.
@@
The existence of two quantum states for a qubit is necessary. One commonly seen example is in trapped ion or neutral atom systems, where the quantum levels of the ions and atoms are used to encode quantum information. Another example is to use the presence or absence of a particle (photons or electrons, etc.) in a single mode as the two quantum states of the qubits. We will discuss this in more detail in Sec. 4.2.3. There may be more than two quantum states in a certain physical system [10]. Other than simply picking two distinct states from all available quantum states, the quantum states need to be controllable (see the following discussion about Operation).
@@
However, some physical platforms may not have sufficient quantum states that fulfill even the minimum requirements of a single qubit. The states spanning the state space should be addressable and controllable. When a single physical system is not sufficient to provide such a controllable state space, however, the composite system of multiple physical systems can be utilized to form a qubit. The composite system can consist of multiple copies of the same physical systems, e.g., the presence or absence of a single electron in a quantum dot cannot be encoded as qubit states due to the lack of coherent manipulation techniques in the Hilbert space spanned by these two states [8]. In contrast, a different physical system can be utilized to provide the necessary qubit operations. e.g., the presence or absence of a microwave photon can be encoded as a qubit with the help of a superconducting qubit [31, 32, 33, 34, 35].
@@
We note that this requirement is weaker than the requirements of single-qubit completely controllable (Rule 2) [29, 30]. Despite incomplete control over the physical system, if a set of computational universal gates can be executed, and the qubit state can be manipulated to a specific state (e.g., |$\ket{0}$), the construction of a qubit is possible. The arbitrary state of the qubit Hilbert space can be approximately synthesized with the extra cost of quantum gate operations (see the discussion of Operations).
@@
The quantum information encoded in a qubits Hilbert space can be transmitted via various physical systems. In measurement-based quantum computing (MBQC) [14, 15, 36, 17], for instance, information is "teleported" to connected physical qubits in resource states, with the physical qubits being destroyed during computation, especially in photonic systems.
@@
b. Relaxing the operation requirement: The qubit should support a set of computational universal gates.
@@
It is quite costly to require any unitary matrices on an arbitrary number of qubits to be directly implementable as different operations. What is the minimum requirement on the set of supported operations to allow us to at least approximate a given unitary matrix on a number of qubits? This question motivates the research surrounding universal gate sets for QC. It turns out that a discrete set of gates (unitary matrices) on a small number of qubits is enough ($n\leq 3$), e.g., the set of Hadamard gate, CNOT gate, phase gate, and $\pi/8$ gate, is universal [9].
@@
However, relaxing this requirement also results in extra costs. Gate decomposition which decomposes the required unitary matrix into the directly supported gate operations in the gate set can be rather costly. The gate decomposition requires an increasing number of quantum gates applied to the qubits. For example, the Solovay-Kitaev theorem showed that approximating an arbitrary unitary matrix in $SU(d)$ using a discrete set of gates that generate a dense subgroup of $SU(d)$ requires $O(log^{c}(1/\epsilon))$ gates, where $c$ is a constant and $\epsilon$ is the accuracy [9, 37]. The second aspect is that the different universal gate sets supported by the respective physical system can cost extra quantum gates and qubits for the same computation, as highlighted by the difference between ‘computational universal’ and ‘strictly universal’ [38]. One example is the gate set that consists of the Toffoli gate and the Hadamard gate. This gate set is universal [38, 39], but a general unitary on a $n$-qubit system cannot be decomposed with arbitrary accuracy using just Toffoli and Hadamard gates. Instead, the computation process of that unitary operation can be simulated within an arbitrarily small error [38]. However, it costs polynomially many more qubits and gates. Similar discussions about universality can be found in qudit systems [40, 24]. We stress that the cost of decomposing target unitaries on the qubit system should be manageable, e.g., not scaling exponentially more time and physical qubits.
@@
Nevertheless, there are other QC models other than the quantum circuit models. For example, in the MBQC scheme [14, 15, 36, 17], universal computing is driven by the measurements on a prepared entangled state. In this model, the computation can still be mapped back to the circuit model, where the flow of quantum information can be treated as qubits that carry quantum information. The physical systems that support MBQC can still be treated as supporting full universal control, though indirectly with the help of measurements and pre-existing quantum entanglement.
c. Relaxing the connectivity requirement: The qubits need to be connected.
@@
For a QC system supporting a universal gate set applicable up to $n$ qubits, a prerequisite for using these qubits in general QC algorithms is that any subset of $n$ qubits must be capable of performing gates from the universal set. Note that all-to-all connectivity is not necessary. For example, if the universal gate set contains CNOT, Hadamard, Phase, and $\pi/8$ gate [9], this is equivalent to requiring that all qubits can perform the Hadamard, phase and $\pi/8$ gates, and any two qubits can find a path in their connection graph. Even if two qubits are not directly connected, a CNOT gate between them can be implemented using SWAP gates along the path, decomposable into three CNOTs. This less stringent connectivity requirement incurs additional costs compared to all-to-all connectivity.
@@
In the new paradigm of distributed QC models, the “connection” between the two qubits we discussed above is not necessarily a direct coupling between two physical qubits. Instead, quantum communication between two physically remote qubits enables remote gate operations [41, 42], which relies on a reliable channel to transmit flying qubits from one end to the other, or a reliable source of entangled resource states (like Bell states) to be used for gate teleportation [43, 44, 45, 46, 9]. This should also comply with DiVincezo’s criteria for quantum communication (criteria 6 and 7, see Sec. 2.2).
@@
d. Relaxing the coherence requirements: The qubit should have a finite but long enough coherence time.
@@
Quantum coherence is critical for quantum computing. However, physical systems can hardly be completely isolated. The interaction with the external environment causes decoherence of the quantum systems. The coherence time gives a time scale for preserving the quantum coherence. These unwanted interactions also introduce decoherence errors to the quantum gates applied to the systems. Therefore, the coherence time should be long enough for the gate operations time to perform a meaningful QC on a physical system. This motivates the development of noisy intermediate-scale quantum (NISQ) algorithms, e.g., variational quantum algorithms [47].
@@
Another strategy is to suppress noise and extend the coherence time of the logical qubits by quantum error correction (QEC) codes [9, 48]. By combining multiple physical qubits into a logical qubit, QEC codes use entangled states as qubit states to detect and correct errors. Increasing the code distance, achieved by using more physical qubits per logical qubit, allows for exponentially suppressing logical error rates as long as noise-induced error rates remain below the code threshold. What’s more, QEC not only extends the coherence time of logical qubits but also suppresses noises from other sources, e.g., imperfect control during gate implementation and cross-talk between adjacent qubits, etc.
@@
e. Relaxing the readout requirement: The qubit should have a way of performing measurements on qubit states.
@@
Furthermore, even the direct projective measurements on the physical systems can be relaxed if we consider heterogeneous QC architectures, where the qubits used in different functional units inside the QC units [49, 50, 51] can be different. The role of the qubit used in the quantum processing units and quantum memory units is to process and store quantum information, which means direct measurements on them are not necessary. We can think of using another physical system, that supports qubit measurement functionalities and can interact with the qubits in the quantum memory modules to extract the stored quantum states. Therefore, when the measurement of the quantum information is needed, it is performed on the second type of qubits by swapping the information out from the computing qubits. A more detailed discussion is in Sec. 4.1.5.
@@
Figure 3: The algorithm-device co-design. Quantum algorithms and applications may only have partial demand on the AQM, while the physical systems can only have partial support of the AQM. The co-design of quantum algorithms and physical systems can be beneficial. While physical systems can only offer partial support for the AQM (the blue arrow), there are quantum algorithms that only partially require the AQM (the red arrow). The AQM provides the necessary tool for quantum co-design, which matches the demand from the quantum algorithms and the support from the physical systems (the gray arrows).
4 AQM for Algorithm-Device codesign
In this section, we discuss the algorithm-device co-design opportunities demonstrated by Fig. 3. Specifically, a quantum application may not require the complete support of the AQM, and hence a physical system that fulfills its requirements can still be useful, even without building a qubit, and vice versa. Therefore, the algorithm-device codesign opportunities by partially breaking the abstraction layer of qubits can be beneficial for the current stage of QC system design. Specifically, in Sec. 4.1, we discuss the algorithm and applications that can be supported without the complete AQM, while in Sec. 4.2, we discuss different physical systems supporting the AQM or quantum applications. In Sec. 4.3, we discuss co-design opportunities of the physical systems.